OpenStudy (anonymous):

A waffle cone has a height of 7 inches and a diameter of 3 inches. What is the volume of ice cream that can be contained within the cone? Use 3.14 for pi. Round your answer to the nearest hundredth. 15.43 cubic inches 16.49 cubic inches 17.86 cubic inches 18.89 cubic inches

OpenStudy (akashdeepdeb):

Volume of a cone is given by 1/3 * pi * r * r * h r = d/2 r = radius d = diameter h = height

OpenStudy (anonymous):

yes i understand that but i dont get 1/3

OpenStudy (automaticloveletter):

I think it's 1/3 because there's three sides in a triangle and a cone is just a 3d form of a triangle. The formula is what @AkashdeepDeb gave you so all you have to do is plug in the numbers in your calculator (or however you get your answers) and you'll get it. To find the radius when given the diameter is simple (just divide by two) Explained Answer: 1/3 * 3.14 * 1.5 * 1.5 * 7 = 16.485 Which can be rounded to --> 16.49 <--

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

well i got a 80 on my test so what ever it was was right

OpenStudy (automaticloveletter):

Whoops, why is it 1/3 then?

OpenStudy (automaticloveletter):

Well either way the answer was right but I'm not sure why it's 1/3 sorry about that lol!

OpenStudy (akashdeepdeb):

 https://proofwiki.org/wiki/Volume_of_Right_Circular_Cone

OpenStudy (akashdeepdeb):

It's okay. :) You'll probably understand the proof a little later.